OpenStudy (anonymous):

the sum of 4 consecutive whole numbers is 54 what are the 4 numbers?

OpenStudy (anonymous):

let the first number be x. so consecutive numbers are x, x+1, x+2,x+3 sum of those 4 numbers = 4x+6 which equals 54 4x +6 =54 4x = 48 x =12. so numbers are 12,13,14,15

jimthompson5910 (jim_thompson5910):

x+(x+1)+(x+2)+(x+3) = 54 4x+6 = 54 4x = 54-6 4x = 48 x = 48/4 x = 12 ---------------- x = 12 x+1 = 13 x+2 = 14 x+3 = 15 So the four numbers are: 12, 13, 14, 15